OpenStudy (anonymous):

find slope-intercept form slope 5/7, ordered pair (0,9/4)

OpenStudy (anonymous):

plug your known values into y=mx+b and solve

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

you know the coordinates are (x,y) x=0 y=(9/4) and your slope is m=4 plug them in and you get 9/4=(5/7)(0)+b

OpenStudy (anonymous):

i need to solve for y

OpenStudy (anonymous):

i have y-9/4=5/7x-(0) but i dont know how to solve that

OpenStudy (anonymous):

i just solved for the b, now we can plug in all our value for the equation. y=(5/7)x+(9/4)

OpenStudy (anonymous):

that is the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much for you help, i am new to this!

OpenStudy (anonymous):

there is two big steps for this, first you sub in all your known values and solve for what ever else you need (in this case b). after you fine the missing value you finial plug all the values in for your final anwser.

OpenStudy (anonymous):

okay, could i continue asking you questions on here?

OpenStudy (anonymous):

(y)=2(x+4) i need to find the points and the slope?